# On modules whose dual is of finite Gorenstein dimension: The $k$-torsion, homological criteria, and applications in modules of differentials of order $n$

### v4

@@ 8. Applications on the $k$-torsion of the modules of differentials
Let $X$ and $Y$ be locally Noetherian spaces and let $k$ be a non-negative integer, $x\in X$, such that ${\rm Der}^{n}_{{(Y,y)}}(X,x)$ has locally finite Gorenstein dimension on $\widetilde{X}^{k-1}(\mathcal{O}_{X,x})$. Then the following conditions are equivalent:
@@
By Fact 7.4(2), we have $\Omega_{(X,x)/(Y,\varphi(x)}^{(n)}=\Omega_{\mathcal{O}_{X,x}/\mathcal{O}_{Y,\varphi(x)}}^{(n)}$. Therefore, $\Omega_{(X,x)/(Y,\varphi(x))}^{(n)}$ is an $\mathcal{O}_{X,x}$-module. Moreover, we know that ${\rm Der}^{n}_{{(Y,\varphi(x))}}(X,x)=(\Omega_{(X,x)/(Y,\varphi(x))}^{(n)})^{\ast}$. Now, the shows of the equivalences follow from Theorem 3.2. ∎
As an immediate consequence of Theorem 8.1 and Corollaries 3.3 and 3.4, we have the following corollaries.
@@
Let $X$ and $Y$ be locally Noetherian spaces and let $x\in X$, such that ${\rm Der}^{n}_{{(Y,y)}}(X,x)$ has locally finite Gorenstein dimension on ${\rm Ass}_{\mathcal{O}_{X,x}}(\mathcal{O}_{X,x})$. Then the following conditions are equivalent:
@@
Let $X$ and $Y$ be locally Noetherian spaces and let $x\in X$ such that ${\rm Der}^{n}_{{(Y,y)}}(X,x)$ has locally finite Gorenstein dimension on $\widetilde{X}^{1}(\mathcal{O}_{X,x})$. Then the following conditions are equivalent:
@@
Let $\varphi:X\rightarrow Y$ be a morphism of schemes which is locally of finite type. If we consider $X$ and $Y$ as affine Noetherian schemes, that is, $X=\operatorname{Spec}(R)$ and $Y=\operatorname{Spec}(S)$ where $R$ and $S$ are local Noetherian rings. By Fact 7.4(2) and Lemma 7.5(ii), if we consider $x\in X$ and $\varphi(x)\in Y$ closed points respectively, then we have $\Omega_{(X,x)/(Y,\varphi(x))}^{(n)}=\Omega^{(n)}_{R/S}$ and ${\rm Der}^{n}_{{(Y,\varphi(x))}}(X,x)={\rm Der}^{n}_{S}(R)$. Moreover, these are finitely generated $R$-modules.
@@
From Remark 8.4, Theorem 8.1, and Corollaries 8.2 and 8.3, we obtain similar results in the affine context, that is, the derivation modules and Kähler differential modules defined over algebras.
Some application in algebraic variety and analytic variety
Before establishing our significant results that connect algebraic geometry and analytic geometry, let’s recall the following definitions: Let $R$ be a (Noetherian) ring, $M$ be an (finitely generated) $R$-module and let $f_{M}:M\to M^{\ast\ast}$ be an $R$-homomorphism. We say that $M$ is torsionless, if the map $f_{M}$ is injective, and $M$ is reflexive if $f_{M}$ is an isomorphism. If $f_{M}$ fails to be surjective, we say that $M$ has cotorsion and we denote the cotorsion by $\text{cotor }M:=\text{coker }(f_{M})$. Note that, by Definition 2.1(5), torsionless is equivalent to 1-torsionfree, and reflexive is equivalent to 2-torsionfree ([50, Proposition 5]).
The torsion submodule of $M$ is defined as the kernel of the natural map $\theta:M\rightarrow M\otimes_{R}{Q(R)}$ where $Q(R)$ is the total quotient ring of $R$. One says that $M$ is torsion-free if $\text{tor}M=0$, and $M$ is a torsion module if $\text{tor}M=M$.
8.5 Remark.
Clearly the torsion submodule $\text{tor}M$ is contained in ker $\text{ker }f_{M}$. If $R$ is a Noetherian domain and $M$ is finitely generated, one can show that ker $\text{ker }f_{M}=\text{tor}M$. Thus, the concepts of torsionless and torsion-free are equivalent ([23, p.70]). If $R$ is only a reduced, Noetherian ring and $M$ is finitely generated, still the concepts are equivalent ([45]). Therefore, in the same situations as above, torsionless, torsion-free, and 1-torsionfree are equivalent.
We shall apply these notions to $(X,\mathcal{O}_{X})$, ringed spaces locally Noetherian (or analytic space), and the stalk at a point $x\in X$ of a coherent sheaf $\mathcal{F}$ viewed as an $\mathcal{O}_{X,x}$-module. For a coherent sheaf $\mathcal{F}$, define the torsion and cotorsion sheaves, tor $\text{tor }\mathcal{F}$ and cotor $\text{cotor }\mathcal{F}$, as the kernel and cokernel sheaves of the natural map $f_{\mathcal{F}}:\mathcal{F}\rightarrow\mathcal{F}^{**}$. At the stalk level, this gives an exact sequence of $\mathcal{O}_{X,x}$-modules
0tor $0\rightarrow\text{tor }\mathcal{F}_{x}\rightarrow\mathcal{F}_{x}\rightarrow(\mathcal{F}^{**})_{x}\rightarrow\text{cotor }\mathcal{F}_{x}\rightarrow 0,$
where the middle map is the evaluation map $f_{{\mathcal{F}}_{x}}$ upon the identification $(\mathcal{F}^{**})_{x}\cong(\mathcal{F}_{x})^{**}$.
8.6 Definition.
Let $X$ be a scheme(or analytic space), and let $x\in X$. Then $\mathfrak{m}_{X,x}/\mathfrak{m}_{X,x}^{2}$ is a vector space over $k(x):=\mathcal{O}_{X,x}/\mathfrak{m}_{X,x}$, called Zariski cotangent space to $X$ at $x$. The Zariski tangent space of $X$ in $x$ is, by definition, the dual vector space
$T_{x}X:=(\mathfrak{m}_{X,x}/\mathfrak{m}_{X,x}^{2})^{\vee}.$
Recall that a locally Noetherian $X$ is said to be nonsingular (or regular) at $x\in X$ if the Zariski tangent space to $X$ at $x$ has dimension equal to $\dim(X,x)$, that is, $\mathcal{O}_{X,x}$ is regular; otherwise, we say that $X$ is singular at $x$. We say that $X$ is regular (or smooth) if it is regular at all of its points. Denote by $\operatorname{Reg}(X)$ and $\operatorname{Sing}(X)$ the regular locus and the singular locus of $X$, respectively. For $x\in X$, the codimension of the singular locus of $X$ at $x$ is defined as $\text{codim}_{x}\text{Sing}\,X:=\text{dim}_{x}X-\text{dim}_{x}\text{Sing}\,X$.
We say that $X$ has isolated singularity if $\operatorname{Sing}(X)$ has dimension at most zero (i.e., equal to 00 or $\infty$- ∞).
8.7 Definition (Cohen-Macaulay Sheaf).
@@
A coherent sheaf $\mathcal{F}$ over a scheme $X$ is said to be (maximal) Cohen-Macaulay if, for every point $x\in X$, the stalk $\mathcal{F}_{x}$ is a (maximal) Cohen-Macaulay module over the local ring $\mathcal{O}_{X,x}$. $X$ is a complete intersection at $x\in X$, if the stalk $\mathcal{O}_{X,x}$ is complete intersection ring. A point $x$ in $X$ is said to be normal if $\mathcal{O}_{X,x}$ is a normal local ring.
In the last part, we considered either one of the following two possible settings, one analytic and one algebraic. More specifically, we consider a reduced complex analytic variety $X$ with the structure sheaf $\mathcal{O}_{X}$, such that, at $x\in X$, the stalk $\mathcal{O}_{X,x}$ is isomorphic to a reduced local $\mathbb{C}$-algebra, which is a quotient of a convergent power series ring $\mathbb{C}\{z_{1},\dots,z_{n}\}$. We also consider an algebraic variety $X$ over an algebraically closed field $k$ of characteristic 00 with the structure sheaf $\mathcal{O}_{X}$, such that, at $x\in X$, the stalk $\mathcal{O}_{X,x}$ is isomorphic as a $k$-algebra, reduced to a quotient of a regular local ring by an ideal.
Recent results on the connection between torsion and cotorsion and reflexivity of Kähler differential modules on algebraic and analytic varieties can be found in the papers by Graf and Miller-Vassiliadou ([22, 39]). Next, we establish a connection between the results of Graf and Miller-Vassiliadou with $k$-torsionfree, $k$-syzygy, and $(\widetilde{S}_{k})$ where $k=1,2$.
@@
Let $X$ be either a complex analytic variety or an algebraic variety over an algebraically closed field of characteristic zero. Assume that $X$ is a complete intersection at a normal point $x\in X$ of dimension $\dim_{x}X=d>0$. Suppose $x$ is a singular point of $X$. Then, for $1\leq n\leq d$, the following conditions are equivalent:
  1. (1)
    $\operatorname{codim}_{x}(\operatorname{Sing}(X))\leq n\leq d.$
  2. (2)
    $\Omega^{(n)}_{X,x}$ is not $1$-torsionfree.
  3. (3)
    $\Omega^{(n)}_{X,x}$ is not $1$-syzygy.
  4. (4)
    $\Omega^{(n)}_{X,x}$ does not satisfy $(\widetilde{S}_{1})$.
@@
The equivalences $(2)\Leftrightarrow(3)\Leftrightarrow(4)$ follow from Corollary 8.2. Since $X$ is reduced, by hypothesis, moreover, in this case, as noted in Remark 8.5, $1$-torsionfree and torsion-free are equivalent. Thus, the equivalence $(2)\Leftrightarrow(1)$ follows from [39, Corollary 3.1]. ∎
8.9 Proposition.
Let $X$ be either a complex analytic variety or an algebraic variety over an algebraically closed field of characteristic zero. Assume that $X$ is a complete intersection at a normal point $x\in X$ of dimension $\dim_{x}X=d>0$. Suppose $x$ is a singular point of $X$. Then, for $1\leq n\leq d$, the following conditions are equivalent:
  1. (1)
    $\operatorname{codim}_{x}(\operatorname{Sing}(X))-1\leq n\leq d$.
  2. (2)
    $\Omega^{(n)}_{X,x}$ is not $2$-torsionfree.
  3. (3)
    $\Omega^{(n)}_{X,x}$ is not $2$-syzygy.
  4. (4)
    $\Omega^{(n)}_{X,x}$ does not satisfy $(\widetilde{S}_{2})$.
  5. (5)
    $\Omega^{(n)}_{X,x}$ is not reflexive.
Proof.
The equivalences $(2)\Leftrightarrow(3)\Leftrightarrow(4)\Leftrightarrow(5)$ follow from Corollary 8.3.
Since reflexive is equivalent to 2-torsionfree ([50, Proposition 5]) and $X$ is reduced, the equivalence $(5)\Leftrightarrow(1)$ follows from [39, Corollary 3.1]. ∎
9. Applications on Herzog-Vasconcelos’s Conjecture
One of the main conjectures, which may have motivated the study of properties of the Kähler differential modules and the regularity of schemes or analytic spaces, is the famous Lipman-Zariski conjecture [33], which states the following: Let $X$ be a complex variety such that the tangent sheaf $\mathcal{T}_{X}:=\mathcal{H}om_{\mathcal{O}_{X}}(\Omega_{X}^{(1)},\mathcal{O}_{X})$ is locally free. Then $X$ is smooth. This conjecture remains widely open. Following the same direction as the Zariski-Lipman Conjecture, there is also a homological version proposed by Herzog and Vasconcelos as:
9.1 Conjecture (Herzog-Vasconcelos’s conjecture).
Let $R$ be a local Noetherian ring. If ${\rm pd}_{R}{\rm Der}_{k}(R)<\infty$, then ${\rm Der}_{k}(R)$ is a free $R$-module.
9.2 Conjecture (Strong Zariski-Lipman Conjecture).
Let $R$ be a local Noetherian ring. If ${\rm pd}_{R}{\rm Der}_{k}(R)<\infty$, then $R$ is regular.
These conjectures, particularly the Lipman-Zariski conjecture when $n=1$, have been extensively studied and resolved in special cases. References of Lipman-Zariski Conjecture can be found in Miller-Vassiliadou [39, Section 4]. While progress has been made for $n>1$, it has been to a lesser extent compared to $n=1$. Further investigations can be found in Graf [22] and Miller-Vassiliadou [39, Section 4]. It’s worth noting that these conjectures are generally false for positive characteristic. For instance, for $n=1$, consider the surface $X\subset\mathbb{A}^{3}_{k}$ over a perfect field $k$ of characteristic $p$ defined by the equation $xy-z^{n}=0$, where $p$ divides $n$. Then one can see that $\mathcal{T}_{X}$ is free (see [33, p. 892]). Example for $n>1$, see [35]. Due to this, from now on our spaces or rings will be equicharacteristic and of characteristic zero.
Next, we present conjectures which are generalizations of 9.1 and 9.2, where SGHVC was introduced by Ludington [35] and GHVC by Graf [22] in the analytical context. Le $M$ be an $R$-module. Since, $\operatorname{pd}_{R}(M)=\sup\{\operatorname{pd}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})|\mathfrak{p}\in\operatorname{Spec}(R)\}.$ Then, the projective dimension of a coherent sheaf over $X$ is defined as
$\operatorname{pd}_{\mathcal{O}_{X}}\mathfrak{F}={\rm sup}\{\operatorname{pd}_{\mathcal{O}_{X,x}}{\mathcal{F}_{x}}|x\in X\}.$
9.3 Questions.
  • (i)
    (Generalizations of Herzog-Vasconcelos’s conjecture (GHVC)) For some integer $n\geq 1$, ${\rm pd}_{\mathcal{O}_{X}}{\mathcal{D}er}_{Y}^{n}(X)\,<\,\infty$. Under what assumptions on $X$ and $Y$, and for which values of $n$, does this imply that ${\mathcal{D}er}_{Y}^{n}(X)$ is locally free?
  • (i)
    (Strong generalizations of Herzog-Vasconcelos’s conjecture (SGHVC)) For some integer $n\geq 1$, ${\rm pd}_{\mathcal{O}_{X}}{\mathcal{D}er}_{Y}^{n}(X)\,<\,\infty$. Under what assumptions on $X$ and $Y$, and for which values of $n$, does this imply that $X$ is locally smooth?
  • (iii)
    In particular, consider $X=\operatorname{Spec}(R)$ and $Y=\operatorname{Spec}(k)$ as in Remark 8.4. Assume that for some integer $n\geq 1$, ${\rm pd}_{R}\operatorname{Der}_{k}^{n}(R)<\infty$. Under what assumptions on $R$, and for which values of $n$ does this imply that $\operatorname{Der}_{k}^{n}(R)$ is free (resp. smooth)?
The following results provide a partial answer to the SGHV conjecture when $X$ is Cohen-Macaulay and the Kähler differential module $\Omega_{X/Y}^{(n)}$ is Cohen-Macaulay or maximal Cohen-Macaulay.
9.4 Proposition.
Let $X$ and $Y$ be locally Noetherian spaces, with $X$ being Cohen-Macaulay, and suppose that ${\rm pd}_{\mathcal{O}_{X}}{\mathcal{D}er}_{Y}^{n}(X)\,<\,\infty$ for some $n\geq 1$. Then ${\mathcal{D}er}_{Y}^{n}(X)$ is locally free if one of the following conditions holds:
  • (1)
    If $\Omega_{X/Y}^{(n)}$ is maximal Cohen-Macaulay.
  • (2)
    If $\Omega_{X/Y}^{(n)}$ is Cohen-Macaulay.
Proof.
@@ 9. Applications on Herzog-Vasconcelos’s Conjecture
It is suffices to show it at the level of stalks. Note that, ${\rm pd}_{\mathcal{O}_{X}}{\mathcal{D}er}_{Y}^{n}(X)\,<\,\infty$ if and only if $\operatorname{pd}_{\mathcal{O}_{X,x}}\left({\rm Der}^{n}_{\mathcal{O}_{Y,\varphi(x)}}(\mathcal{O}_{X,x})\right)<\infty$ for all $x\in X$. Moreover, $\Omega_{X/Y}^{(n)}$ is Cohen-Macaulay (maximal) if and only if $\Omega_{(X,x)/(Y,y)}^{(n)}$ is Cohen-Macaulay (maximal) for all $x\in X$.
Now, by Corollary 5.2 or Corollary 5.3, $\Omega_{(X,x)/(Y,y)}^{(n)}$ is an $\mathcal{O}_{X,x}$-free module for all $x\in X$. Thus, by the universal property, we obtain that ${\rm Der}^{n}_{{(Y,y)}}(X,x)$ is free for all $x\in X$. ∎
Note that in Proposition 9.4 for $n=1$, if we consider $X=\operatorname{Spec}(R)$, $Y=\operatorname{Spec}(k)$ as in Remark 8.4, the assumption ${\rm pd}_{R}({\rm Der}_{k}(R))<\infty$ implies that $R$ is Gorenstein (see [25]). Thus, by Theorem 4.2 and Corollary 4.4, ${\rm Der}_{k}(R)$ is totally reflexive, hence ${\rm Der}_{k}(R)$ is free.
The following results provide a partial answer to the SGHV conjecture for complete intersections.
9.5 Proposition.
Let $X$ be either a complex analytic variety or an algebraic variety over an algebraically closed field of characteristic zero. Assume that $X$ is a complete intersection at a point $x\in X$ of dimension $\text{dim}_{x}X=d>0$ and $\Omega_{X,x}^{(n)}$ is (maximal) Cohen-Macaulay for some $11\leq n\leq d-$ 1. Let the codimension of the singular locus of $X$ at $x$ be at least three. If $\operatorname{pd}_{\mathcal{O}_{X,x}}\left({\rm Der}^{n}_{k}(\mathcal{O}_{X,x})\right)<\infty$ at $x$, then $x$ is a smooth point of $X$.
Proof.
In the proof of Theorem 9.4, it is noted that such a result holds at the level of stalks. Additionally, in the same proof or by Corollary 5.2, Corollary 5.3, we show that $\Omega_{X,x}^{(n)}$ is free as an $\mathcal{O}_{X,x}$-module. Therefore, $\Omega_{X,x}^{[n]}$ is a $\mathcal{O}_{X,x}$-free module, because $\Omega_{X,x}^{[n]}=\left(\Omega_{X,x}^{(n)}\right)^{\ast\ast}$. Now by hypothesis and from [39, Theorem 4.1], we get the result. ∎

### v5

@@ 8.1. Applications on the $k$-torsion of the modules of differentials
Let $X$ and $Y$ be locally Noetherian schemes, $x\in X$ and $k$ be a non-negative integer such that ${\rm Der}^{n}_{{(Y,y)}}(X,x)$ has locally finite Gorenstein dimension on $\widetilde{X}^{k-1}(\mathcal{O}_{X,x})$. Then the following conditions are equivalent:
@@
By Facts 7.4(2), we have $\Omega_{(X,x)/(Y,y)}^{(n)}=\Omega_{\mathcal{O}_{X,x}/\mathcal{O}_{Y,y}}^{(n)}$. Therefore, $\Omega_{(X,x)/(Y,y)}^{(n)}$ is an $\mathcal{O}_{X,x}$-module. Moreover, we know that ${\rm Der}^{n}_{{(Y,y)}}(X,x)=\left(\Omega_{(X,x)/(Y,y)}^{(n)}\right)^{\ast}$. Now, the required equivalences follow from Theorem 3.2. ∎
As an immediate consequence of Theorem 8.1, or Corollaries 3.3 and 3.4, we have the following corollaries.
@@
Let $X$ and $Y$ be locally Noetherian schemes, and let $x\in X$ such that ${\rm Der}^{n}_{{(Y,y)}}(X,x)$ has locally finite Gorenstein dimension on ${\rm Ass}_{\mathcal{O}_{X,x}}(\mathcal{O}_{X,x})$. Then the following conditions are equivalent:
@@
Let $X$ and $Y$ be locally Noetherian schemes, and let $x\in X$ such that ${\rm Der}^{n}_{{(Y,y)}}(X,x)$ has locally finite Gorenstein dimension on $\widetilde{X}^{1}(\mathcal{O}_{X,x})$. Then the following conditions are equivalent:
@@
Let $\varphi:X\rightarrow Y$ be a morphism of schemes which is locally of finite type. If we consider $X$ and $Y$ as affine Noetherian schemes, that is, $X=\operatorname{Spec}(R)$ and $Y=\operatorname{Spec}(S)$ where $R$ and $S$ are local Noetherian rings. By Facts 7.4(2) and Lemma 7.5(ii), if we consider $x\in X$ and $\varphi(x)\in Y$ closed points respectively, then we have $\Omega_{(X,x)/(Y,\varphi(x))}^{(n)}=\Omega^{(n)}_{R/S}$ and ${\rm Der}^{n}_{{(Y,\varphi(x))}}(X,x)={\rm Der}^{n}_{S}(R)$. Moreover, these are finitely generated $R$-modules.
@@
From Remark 8.4, Theorem 8.1, and Corollaries 8.2 and 8.3, we can obtain similar results in the affine context, that is, the derivation modules and Kähler differential modules defined over algebras.
8.2. Applications on Herzog-Vasconcelos’s conjecture
One of the main conjectures, which may have motivated the study of properties of the Kähler differential modules and the regularity of schemes, is the famous Lipman-Zariski conjecture [Lipman], which states the following: Let $X$ be a complex variety such that the tangent sheaf $\mathcal{T}_{X}:=\mathcal{H}om_{\mathcal{O}_{X}}\left(\Omega_{X}^{(1)},\mathcal{O}_{X}\right)$ is locally free. Then $X$ is smooth. This conjecture remains widely open. Lipman-Zariski conjecture, have been extensively studied and resolved in special cases. References of Lipman-Zariski conjecture can be found in Miller-Vassiliadou [37, Section 4]. It is important to note that this conjecture is generally false for positive characteristic. For example, consider the surface $X\subset\mathbb{A}^{3}_{k}$ over a perfect field $k$ of characteristic $p$, defined by the equation $xy-z^{n}=0$, where $p$ divides $n$. Then one can see that $\mathcal{T}_{X}$ is free (see [Lipman, p. 892]). Following the same direction as the Zariski-Lipman conjecture, there is also a homological version proposed by Herzog and Vasconcelos:
8.5 Conjecture (Herzog-Vasconcelos’s conjecture).
Let $R$ be a local $k$-algebra, where $k$ is a field of characteristic zero. If ${\rm pd}_{R}({\rm Der}_{k}(R))<\infty$, then ${\rm Der}_{k}(R)$ is a free $R$-module.
This conjecture, unlike the former, appears to be widely open (with the exception of a few specific cases; see, e.g., [He-M, Section 4] and [HV, PerezNeto]).
Motivated by Herzog-Vasconcelos’s conjecture and the definitions from the previous sections, we can present the following question. But first, let us define: Let $M$ be an $R$-module. Since $\operatorname{pd}_{R}(M)=\sup\{\operatorname{pd}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\mid\mathfrak{p}\in\operatorname{Spec}(R)\},$ the projective dimension of a coherent sheaf $\mathcal{F}$ over an scheme $X$ is defined as $\operatorname{pd}_{\mathcal{O}_{X}}\mathcal{F}=\sup\{\operatorname{pd}_{\mathcal{O}_{X,x}}(\mathcal{F}_{x})\mid x\in X\}.$
8.6 Questions.
(Generalization of Herzog-Vasconcelos’s conjecture (GHVC)).
  • (i)
    Let $n\geq 1$ and ${\rm pd}_{\mathcal{O}_{X}}{\mathcal{D}er}_{Y}^{n}(X)\,<\,\infty$. Under what assumptions on $X$ and $Y$, and for which values of $n$, does this imply that ${\mathcal{D}er}_{Y}^{n}(X)$ is locally free?
  • (ii)
    In particular, consider $X=\operatorname{Spec}(R)$ and $Y=\operatorname{Spec}(S)$ as in Remark 8.4. Assume that for some integer $n\geq 1$, ${\rm pd}_{R}(\operatorname{Der}_{S}^{n}(R))<\infty$. Under what assumptions on $R$, and for which values of $n$ does this imply that $\operatorname{Der}_{S}^{n}(R)$ is free?
In the context of Question 8.6(ii), a generalized form of the Zariski-Lipman Conjecture can be stated as follows: If $k$ is a field and $R$ is a $k$-algebra, then $R$ is a regular local ring if $\operatorname{Der}_{k}^{n}(R)$ is a free $R$-module for some $n\geq 1$. It should also be noted that for $n>1$, Question 8.6(ii) generally fails when $S=k$ is a field of positive characteristic. For a counterexample, see [Lu].
@@ 8.2. Applications on Herzog-Vasconcelos’s conjecture
A coherent sheaf $\mathcal{F}$ over a locally Noetherian scheme $X$ is said to be Cohen-Macaulay if, for every point $x\in X$, the stalk $\mathcal{F}_{x}$ is a Cohen-Macaulay module over the local ring $\mathcal{O}_{X,x}$.
The following result provides a partial answer to the GHVC conjecture when $X$ and $Y$ are locally Noetherian schemes with $X$ and the Kähler differential module $\Omega_{X/Y}^{(n)}$ being Cohen-Macaulay.
@@
Let $X$ and $Y$ be locally Noetherian schemes, with $X$ Cohen-Macaulay, and suppose that ${\rm pd}_{\mathcal{O}_{X}}{\mathcal{D}er}_{Y}^{n}(X)<\infty$ for some $n\geq 1$. If $\Omega_{X/Y}^{(n)}$ is Cohen-Macaulay, then ${\mathcal{D}er}_{Y}^{n}(X)$ is locally free.
@@
It is suffices to show it at the level of stalks. Note that, ${\rm pd}_{\mathcal{O}_{X}}{\mathcal{D}er}_{Y}^{n}(X)<\infty$ if and only if $\operatorname{pd}_{\mathcal{O}_{X,x}}\left({\rm Der}^{n}_{\mathcal{O}_{Y,y}}(\mathcal{O}_{X,x})\right)<\infty$ for all $x\in X$. Moreover, $\Omega_{X/Y}^{(n)}$ is Cohen-Macaulay if and only if $\Omega_{(X,x)/(Y,y)}^{(n)}$ is Cohen-Macaulay for all $x\in X$. Now, by Corollary 5.3, $\Omega_{(X,x)/(Y,y)}^{(n)}$ is a free $\mathcal{O}_{X,x}$-module for all $x\in X$. Thus, by the universal property, we obtain that ${\rm Der}^{n}_{{(Y,y)}}(X,x)$ is free for all $x\in X$. ∎